OpenStudy (anonymous):

Find and simplify the difference quotient [ f(x+h) - f(x) ] / h, h does not equal 0 f(x) = x^2

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

answer is not 2x and i need an explanation

OpenStudy (anonymous):

Ohh, sorry I thought it was when h tends to 0 xD

OpenStudy (anonymous):

oh haha idk how to do it im confused on how to get the 2x

OpenStudy (aum):

f(x) = x^2 f(x+h) = (x+h)^2 [ f(x+h) - f(x) ] / h = [ (x+h)^2 - x^2 ] / h = (x+h+x)(x+h-x) / h = (2x+h)(h) / h = (2x+h)

OpenStudy (aum):

As h->0, 2x+h -> 2x

OpenStudy (anonymous):

wait how does (x+h)^2 - x^2 go to (x+h+x)(x+h-x)

OpenStudy (aum):

There is an algebraic identity for the difference of two squares: \(a^2 - b^2 = (a+b)(a-b) \) Therefore, \((x+h)^2 - x^2 = (x+h+x)(x+h-h) = (2x+h)(x)\)

OpenStudy (anonymous):

OHHHH okay

OpenStudy (aum):

I meant \((x+h)^2 - x^2 = (x+h+x)(x+h-x) = (2x+h)(h)\)

OpenStudy (anonymous):

so how do you get the ending x? i understand (x+h)(x+h) then idk the -x^2

OpenStudy (aum):

typo in my reply before the last one which I corrected in the previous reply.

OpenStudy (anonymous):

im reading the most recent one but i still dont understand the ending x?

OpenStudy (aum):

There is a 'h' at the end not 'x'

OpenStudy (anonymous):

in the second part

OpenStudy (aum):

a^2 - b^2 = (a+b)(a-b) (x+h)^2 - x^2 = ? a = x+h, b = x a + b = x + h + x = 2x + h a - b = x + h - x = h (a+b)(a-b) = (2x+h)(h)

OpenStudy (anonymous):

okay i understand now, thank you so much

OpenStudy (aum):

You are welcome. You can also do this the long way: (x+h)^2 = x^2 + 2xh + h^2 (x+h)^2 - x^2 = x^2 + 2xh + h^2 - x^2 = 2xh + h^2 = (2x+h)(h)